OpenStudy (anonymous):

if each edge of a cube is doubled in length,then the volume of the cube is is multiplied by a factor of??

OpenStudy (amistre64):

try it out

OpenStudy (amistre64):

s*s*s = s^3 2s*2s*2s = 8s^3